# Packing subgraphs in regular graphs

### v1

@@ 6 $H$-packings in dense regular graphs
For all $i\in[r]$, by (B2), $G^{\prime}[Z_{i}]$ has average degree at least $d-\eta|Z_{i}|$, by (B3), $G^{\prime}[Z_{i}]$ has no $\zeta$-sparse cuts and, by (B4), either $G^{\prime}[Z_{i}]$ is bipartite and balanced, or one needs to remove at least $\gamma n^{2}\geq\gamma|Z_{i}|^{2}$ edges to make it bipartite, and $d\geq cn\geq c|Z_{i}|$. Thus, for every $i\in[r]$, we can apply Corollary˜5.2 with $G^{\prime}[Z_{i}]$ playing the role of $G$, to obtain a $K_{t,t}$-packing of $G^{\prime}[Z_{i}]$ covering all but at most 2t12t-1 vertices in $Z_{i}$. Moreover, $G[L]$ has a perfect $K_{t,t}$-packing by (B5), and $|V(G)-(Z_{1}\cup\ldots\cup Z_{r}\cup L)|\leq 64r^{2}\cdot\frac{t}{c}\cdot\left(\frac{e}{8\zeta}\right)^{t}$ by (B1). This yields a $K_{t,t}$-packing of $G$ covering all but at most the following number of vertices.
@@ 7 Packing subdivisions in dense regular graphs
In this section, we prove Theorem˜1.7. We start by giving a detailed sketch of its proof here. Let $F$ be a graph, and let $0\leq c\leq 1$. Let $G$ be a $d$-regular graph of order $n$, where $d\geq cn$. We start, as before, by partitioning the vertex set of $G$ into sets $Z_{1},\ldots,Z_{r}$ such that the subgraphs $G[Z_{i}]$, $i\in[r]$ induced by these sets are expanders with few edges between them (using Lemma 3.2). Let $\{X_{i},Y_{i}\}$ be a bipartition of $G$ maximising the number of edges in $G[X_{i},Y_{i}]$. Let $G^{\prime}$ be the graph obtained from $G$ by removing all edges within $X_{i}$ and $Y_{i}$ for all $i\in[r]$ such that $G[Z_{i}]$ is close to being bipartite. This ensures that for all $i\in[r]$, $G^{\prime}[Z_{i}]$ is either bipartite with the bipartition $\{X_{i},Y_{i}\}$ (with roughly the same average degree and minimum degree as $G[Z_{i}]$) or it is far from being bipartite.
Next, we find short paths $P_{1},\ldots,P_{r}$ (where each $P_{i}$ has both of its ends in $Z_{i}$) whose removal ‘balances’ the expanders that are close to being bipartite. More precisely, for every $i\in[r]$, if $G[Z_{i}]$ is close to being bipartite, then, writing $Q\coloneqq P_{1}\cup\ldots\cup P_{r}$, we have $|X_{i}-V(Q)|=|Y_{i}-V(Q)|$, and $P_{i}$ has one end in $X_{i}$ and the other end in $Y_{i}$. The key tool for building these paths is Lemma˜3.4 due to Gruslys and Letzter [11], which produces a small linear forest $H$ that balances the expanders that are close to being bipartite, while also ensuring that each expander contains zero or two leaves of $H$. To obtain the desired paths $P_{1},\ldots,P_{r}$, we iteratively merge pairs of components of $H$ whose leaves lie in the same expander and, if necessary, we define some paths $P_{i}$ as arbitrary edges within $G^{\prime}[Z_{i}]$.
@@
In the last three steps, we rely on the fact that the expanders $G^{\prime}[Z_{i}]$ are robustly connected via short paths; that is, any two vertices can be joined by a short path that avoids any given small set of forbidden vertices (see Lemma˜7.1). In the final step, we also use the robust Hamiltonicity property of the expanders $G^{\prime}[Z_{i}]$, meaning that for any small set of forbidden vertices $W$ (which is balancing if $G^{\prime}[Z_{i}]$ is bipartite), and any two vertices $x,y\in Z_{i}-W$ (lying in different parts of $G^{\prime}[Z_{i}]$ if it is bipartite), there exists a Hamilton path in $G^{\prime}[Z_{i}]-W$ with ends $x$ and $y$.
The proof of Theorem˜1.7 is given in Section˜7.3, after establishing two preparatory lemmas in the following two subsections.
@@ 7.1 Expanders are robustly connected via short paths
Now we prove (ii). Let $p,p^{\prime}\in V(H^{\prime\prime})$. By (i), $H^{\prime\prime}$ is connected, so it contains a $(p,p^{\prime})$-path. Let $P$ be a shortest $(p,p^{\prime})$-path in $H^{\prime\prime}$. We claim that $P$ has length at most $15/\delta$. Indeed, suppose otherwise, and let $P$ be the path $u_{0}\ldots u_{t}$, where $u_{0}=p$ and $u_{t}=p^{\prime}$ such that $t>15/\delta$. Set $U\coloneqq\{u_{0},u_{3},\ldots,u_{3q}\}$, where $q\coloneqq\lceil 4/\delta\rceil$ (this is well defined because $t>\frac{15}{\delta}\geq 3\cdot(\frac{4}{\delta}+1)\geq 3\cdot\left\lceil\frac{4}{\delta}\right\rceil=3q$). Since $H^{\prime\prime}$ has minimum degree at least $\delta n/4$, the number of edges of $H^{\prime\prime}$ incident to vertices in $U$ is at least $(q+1)\cdot\delta n/4>n$, implying that there is a vertex $v$ in $H^{\prime\prime}$ with at least two neighbours in $U$, say $u_{3i}$ and $u_{3j}$, where $i<j$. Denote by $W$ the walk $u_{0}u_{1}\ldots u_{3i}vu_{3j}\ldots u_{t}$. Note that $W$ is a $(p,p^{\prime})$-walk in $H^{\prime\prime}$. Let $P^{\prime}$ be a $(p,p^{\prime})$-path contained in $W$. Since $W$ is shorter than $P$, it follows that $P^{\prime}$ is also shorter than $P$, a contradiction to the minimality of $P$. This proves (ii), completing the proof of Lemma˜7.1. ∎
@@ 7.2 Balancing the expanders
As further preparation for the proof of Theorem˜1.4, we state the following corollary of Lemma˜3.4, which will allow us to balance the expanders using a suitable collection of paths. This is somewhat similar to Lemma˜4.1, where we also balance the expanders, but using a collection of vertex-disjoint $K_{t,t}$’s instead of paths.
@@
Let $H$ be the linear forest guaranteed by Lemma˜3.4, satisfying properties (P1)(P4). Let $G^{\prime}$ be the subgraph of $G$ obtained by removing all edges with both endpoints in $X_{i}$ or both endpoints in $Y_{i}$, for each $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite.
@@
For $i\in[r]$, suppose that $H_{i-1}$ is already defined such that $|V(H_{i-1})|\leq\xi n+15(i-1)/\delta$; we construct $H_{i}$ as follows. If $H_{i-1}$ has two distinct components, say $P$ and $P^{\prime}$, with leaves $p,p^{\prime}$ in $Z_{i}$, say $p\in V(P)\cap Z_{i}$, $p^{\prime}\in V(P^{\prime})\cap Z_{i}$, let $L_{i}$ be a shortest $(p,p^{\prime})$-path in $G^{\prime}[Z_{i}]-(V(H_{i-1})-\{p,p^{\prime}\})$. Applying (ii) of Lemma˜7.1 with $G[Z_{i}]$, $G^{\prime}[Z_{i}]$, $V(H_{i-1})-\{p,p^{\prime}\}$, $2\xi$ playing the roles of $H$, $H^{\prime}$, $W$, $\xi$, respectively, we obtain that the path $L_{i}$ has length at most $15/\delta$. (Note that Lemma˜7.1 indeed applies because $|V(H_{i-1})|\leq\xi n+15(i-1)/\delta\leq 2\xi n$.) Define $H_{i}\coloneqq H_{i-1}\cup L_{i}$. Note that $|V(H_{i})|\leq|V(H_{i-1})|+15/\delta\leq\xi n+15i/\delta$, as required.
@@
7.3 Proof of Theorem˜1.7
We are now ready to prove Theorem˜1.7.
@@
By applying Lemma˜3.2 to $G$, we obtain positive numbers $\eta,\beta,\gamma,\zeta,\delta$ where
@@
For each $i\in[r]$, let $X_{i},Y_{i}$} be a partition of $Z_{i}$ that maximizes $e_{G}(X_{i},Y_{i})$. Let $G^{\prime}$ be the subgraph of $G$ obtained by removing all edges with both endpoints in $X_{i}$ or both in $Y_{i}$, for each $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite. Let $P_{1},\ldots,P_{r}$ be vertex-disjoint paths satisfying (P’1)(P’3), guaranteed by Lemma˜7.2, and let $Q\coloneqq P_{1}\cup\ldots\cup P_{r}$.
@@ 7.3.1 Constructing a balanced pair of $F$-subdivisions in each expander
Thus, by applying (ii) of Lemma˜7.1 (with $2\xi$ playing the role of $\xi$), we obtain an $(s,s^{\prime})$-path $P$ in $G^{\prime}[Z_{i}]-W$ of length at most $15/\delta$. But then $\mathcal{P}\cup\{P\}$ contradicts the maximality of $\mathcal{P}$. This shows that $|\mathcal{P}|=2e(F)$, as claimed.
Pick a collection $\mathcal{P}$ as guaranteed by the previous paragraph. Take $F_{i}$ to be the subgraph consisting of the paths in $\mathcal{P}$ with ends in $S$, and take $F_{i}^{\prime}$ to be the subgraph consisting of the paths in $\mathcal{P}$ with ends in $S^{\prime}$. We claim that $F_{i}$ and $F_{i}^{\prime}$ satisfy the requirements of Claim˜7.3. Indeed, first note that $F_{i}$ is an $F$-subdivision in $G^{\prime}[Z_{i}]-V(Q)$ whose branch vertices are in $S$, and similarly $F_{i}^{\prime}$ is an $F$-subdivision in $G^{\prime}[Z_{i}]-V(Q)$ whose branch vertices are in $S^{\prime}$. Second, notice that $F_{i}$ and $F_{i}^{\prime}$ are vertex-disjoint, since the paths in $\mathcal{P}$ are internally vertex-disjoint whose internal vertices are in $Z_{i}-(V(Q)\cup S\cup S^{\prime})$. Third, we have $|V(F_{i})|,|V(F_{i}^{\prime})|\leq e(F)\cdot 15/\delta$, since every path in $\mathcal{P}$ has length at most $15/\delta$ by our choice of $\mathcal{P}$. Finally, if $G[Z_{i}]$ is $\beta$-almost-bipartite, then for every path $P\in\mathcal{P}$ with ends in $S$, we have
@@
This proves (22), completing the proof of Claim˜7.3. ∎
@@ 7.3.2 Absorbing the paths $P_{1},\ldots,P_{r}$ and all the uncovered vertices into the subdivisions $F_{i}$
For every $i\in[r]$, let $F_{i},F^{\prime}_{i}$ be the two vertex-disjoint subdivisions of $F$ in $G^{\prime}[Z_{i}]-V(Q)$, as guaranteed by Claim˜7.3. Denote the leaves of $P_{i}$ by $u_{i}$ and $v_{i}$, where $u_{i}\in X_{i}$ and $v_{i}\in Y_{i}$ if $G[Z_{i}]$ is $\beta$-almost-bipartite. Let $x_{i}y_{i}$ be an arbitrary edge of $F_{i}$, where $x_{i}\in X_{i}$ and $y_{i}\in Y_{i}$ if $G[Z_{i}]$ is $\beta$-almost-bipartite. Let $Q_{i}$ be a shortest $(y_{i},u_{i})$-path in $G^{\prime}[Z_{i}]$ whose interior avoids $V(Q)\cup V(F_{i})\cup V(F_{i}^{\prime})$. By (P’1),
@@
Hence, it follows that $Q_{i}$ exists and is of length at most $15/\delta$ by applying (ii) of Lemma˜7.1 (with $\sqrt{\xi}$ playing the role of $\xi$). Let $H\coloneqq G[Z_{i}]$, and let $H^{\prime}\coloneqq G^{\prime}[Z_{i}]$. Then, by Lemma˜7.1, $H^{\prime}$ has no $\zeta/2$-sparse cuts. Now let
@@
showing that $H^{\prime}$ has average degree at least $d-(2\beta/\delta^{2})|V(H^{\prime})|$. Moreover, $H^{\prime}$ is either bipartite or $\gamma$-far-from-bipartite by definition. Therefore, using $|W_{i}|\leq 2\sqrt{\xi}n$, we can apply Lemma˜3.3 with $H^{\prime}$, $W_{i}$, $2\beta/\delta^{2}$, $2\xi$, $\zeta/2$ playing the roles of $G$, $W$, $\eta$, $\xi$, $\zeta$, respectively, to find a Hamilton path $Q_{i}^{\prime}$ in $H^{\prime}-W_{i}=G^{\prime}[Z_{i}]-W_{i}$ with ends $x_{i}$ and $v_{i}$. Note that Lemma˜3.3 indeed applies because when $G[Z_{i}]$ is $\beta$-almost-bipartite (so $H^{\prime}$ is bipartite), (P’3) and (22) together imply that $|X_{i}-W_{i}|=|Y_{i}-W_{i}|$.

### v2

@@ 6 $F$-packings in dense regular graphs
For all $i\in[r]$, by (B2), $G^{\prime}[Z_{i}]$ has average degree at least $d-\eta|Z_{i}|$, by (B3), $G^{\prime}[Z_{i}]$ has no $\zeta$-sparse cuts and, by (B4), either $G^{\prime}[Z_{i}]$ is bipartite and balanced, or one needs to remove at least $\gamma n^{2}\geq\gamma|Z_{i}|^{2}$ edges to make it bipartite, and $d\geq cn\geq c|Z_{i}|$. Thus, for every $i\in[r]$, we can apply Corollary 5.2 with $G^{\prime}[Z_{i}]$ playing the role of $G$, to obtain a $K_{t,t}$-packing of $G^{\prime}[Z_{i}]$ covering all but at most 2t12t-1 vertices in $Z_{i}$. Moreover, $G[L]$ has a perfect $K_{t,t}$-packing by (B5), and $|V(G)-(Z_{1}\cup\ldots\cup Z_{r}\cup L)|\leq 64r^{2}\cdot\frac{t}{c}\cdot\left(\frac{e}{8\zeta}\right)^{t}$ by (B1). This yields a $K_{t,t}$-packing of $G$ covering all but at most the following number of vertices.
@@ 7 Packing subdivisions in dense regular graphs
In this section, we prove Theorem 1.7. We start by giving a detailed sketch of its proof here. Let $F$ be a graph, and let $0\leq c\leq 1$. Let $G$ be a $d$-regular graph of order $n$, where $d\geq cn$. We start, as before, by partitioning the vertex set of $G$ into sets $Z_{1},\ldots,Z_{r}$ such that the subgraphs $G[Z_{i}]$, $i\in[r]$ induced by these sets are expanders with few edges between them (using Lemma 3.2). Let $\{X_{i},Y_{i}\}$ be a bipartition of $G$ maximising the number of edges in $G[X_{i},Y_{i}]$. Let $G^{\prime}$ be the graph obtained from $G$ by removing all edges within $X_{i}$ and $Y_{i}$ for all $i\in[r]$ such that $G[Z_{i}]$ is close to being bipartite. This ensures that for all $i\in[r]$, $G^{\prime}[Z_{i}]$ is either bipartite with the bipartition $\{X_{i},Y_{i}\}$ (with roughly the same average degree and minimum degree as $G[Z_{i}]$) or it is far from being bipartite.
Next, we find short paths $P_{1},\ldots,P_{r}$ (where each $P_{i}$ has both of its ends in $Z_{i}$) whose removal ‘balances’ the expanders that are close to being bipartite. More precisely, for every $i\in[r]$, if $G[Z_{i}]$ is close to being bipartite, then, writing $Q\coloneqq P_{1}\cup\ldots\cup P_{r}$, we have $|X_{i}-V(Q)|=|Y_{i}-V(Q)|$, and $P_{i}$ has one end in $X_{i}$ and the other end in $Y_{i}$. The key tool for building these paths is Lemma 3.4 due to Gruslys and Letzter [11], which produces a small linear forest $H$ that balances the expanders that are close to being bipartite, while also ensuring that each expander contains zero or two leaves of $H$. To obtain the desired paths $P_{1},\ldots,P_{r}$, we iteratively merge pairs of components of $H$ whose leaves lie in the same expander and, if necessary, we define some paths $P_{i}$ as arbitrary edges within $G^{\prime}[Z_{i}]$.
@@
In the last three steps, we rely on the fact that the expanders $G^{\prime}[Z_{i}]$ are robustly connected via short paths; that is, any two vertices can be joined by a short path that avoids any given small set of forbidden vertices (see Lemma 7.1). In the final step, we also use the robust Hamiltonicity property of the expanders $G^{\prime}[Z_{i}]$, meaning that for any small set of forbidden vertices $W$ (which is balancing if $G^{\prime}[Z_{i}]$ is bipartite), and any two vertices $x,y\in Z_{i}-W$ (lying in different parts of $G^{\prime}[Z_{i}]$ if it is bipartite), there exists a Hamilton path in $G^{\prime}[Z_{i}]-W$ with ends $x$ and $y$.
The proof of Theorem 1.7 is given in Section 7.3, after establishing two preparatory lemmas in the following two subsections.
@@ 7.1 Expanders are robustly connected via short paths
Now we prove (ii). Let $p,p^{\prime}\in V(H^{\prime\prime})$. By (i), $H^{\prime\prime}$ is connected, so it contains a $(p,p^{\prime})$-path. Let $P$ be a shortest $(p,p^{\prime})$-path in $H^{\prime\prime}$. We claim that $P$ has length at most $15/\delta$. Indeed, suppose otherwise, and let $P$ be the path $u_{0}\ldots u_{t}$, where $u_{0}=p$ and $u_{t}=p^{\prime}$ such that $t>15/\delta$. Set $U\coloneqq\{u_{0},u_{3},\ldots,u_{3q}\}$, where $q\coloneqq\lceil 4/\delta\rceil$ (this is well defined because $t>\frac{15}{\delta}\geq 3\cdot(\frac{4}{\delta}+1)\geq 3\cdot\left\lceil\frac{4}{\delta}\right\rceil=3q$). Since $H^{\prime\prime}$ has minimum degree at least $\delta n/4$, the number of edges of $H^{\prime\prime}$ incident to vertices in $U$ is at least $(q+1)\cdot\delta n/4>n$, implying that there is a vertex $v$ in $H^{\prime\prime}$ with at least two neighbours in $U$, say $u_{3i}$ and $u_{3j}$, where $i<j$. Denote by $W$ the walk $u_{0}u_{1}\ldots u_{3i}vu_{3j}\ldots u_{t}$. Note that $W$ is a $(p,p^{\prime})$-walk in $H^{\prime\prime}$. Let $P^{\prime}$ be a $(p,p^{\prime})$-path contained in $W$. Since $W$ is shorter than $P$, it follows that $P^{\prime}$ is also shorter than $P$, a contradiction to the minimality of $P$. This proves (ii), completing the proof of Lemma 7.1. ∎
@@ 7.2 Balancing the expanders
As further preparation for the proof of Theorem 1.4, we state the following corollary of Lemma 3.4, which will allow us to balance the expanders using a suitable collection of paths. This is somewhat similar to Lemma 4.1, where we also balance the expanders, but using a collection of vertex-disjoint $K_{t,t}$’s instead of paths.
@@
Let $H$ be the linear forest guaranteed by Lemma 3.4, satisfying properties (P1)(P4). Let $G^{\prime}$ be the subgraph of $G$ obtained by removing all edges with both endpoints in $X_{i}$ or both endpoints in $Y_{i}$, for each $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite.
@@
For $i\in[r]$, suppose that $H_{i-1}$ is already defined such that $|V(H_{i-1})|\leq\xi n+15(i-1)/\delta$; we construct $H_{i}$ as follows. If $H_{i-1}$ has two distinct components, say $P$ and $P^{\prime}$, with leaves $p,p^{\prime}$ in $Z_{i}$, say $p\in V(P)\cap Z_{i}$, $p^{\prime}\in V(P^{\prime})\cap Z_{i}$, let $L_{i}$ be a shortest $(p,p^{\prime})$-path in $G^{\prime}[Z_{i}]-(V(H_{i-1})-\{p,p^{\prime}\})$. Applying (ii) of Lemma 7.1 with $G[Z_{i}]$, $G^{\prime}[Z_{i}]$, $V(H_{i-1})-\{p,p^{\prime}\}$, $2\xi$ playing the roles of $H$, $H^{\prime}$, $W$, $\xi$, respectively, we obtain that the path $L_{i}$ has length at most $15/\delta$. (Note that Lemma 7.1 indeed applies because $|V(H_{i-1})|\leq\xi n+15(i-1)/\delta\leq 2\xi n$.) Define $H_{i}\coloneqq H_{i-1}\cup L_{i}$. Note that $|V(H_{i})|\leq|V(H_{i-1})|+15/\delta\leq\xi n+15i/\delta$, as required.
@@
V(Q)Zi={V(H)Ziif PiHr(V(H)Zi)
V
if PiHr and Li is not defined
(V(H)Zi)
V
(V(Li){p,p})
if PiHr and Li is defined.
@@
denotes the union of the disjoint sets $A$ and $B$.) Hence, (P’3) follows from (P4), (21), and the choice of $P_{i}$ as an edge in $G^{\prime}[Z_{i}]-V(P_{r})$ when $P_{i}\nsubseteq H_{r}$, completing the proof of Lemma 7.2. ∎
7.3 Proof of Theorem 1.7
We are now ready to prove Theorem 1.7.
@@
By applying Lemma 3.2 to $G$, we obtain positive numbers $\eta,\beta,\gamma,\zeta,\delta$ where
@@
For each $i\in[r]$, let $X_{i},Y_{i}$} be a partition of $Z_{i}$ that maximizes $e_{G}(X_{i},Y_{i})$. Let $G^{\prime}$ be the subgraph of $G$ obtained by removing all edges with both endpoints in $X_{i}$ or both in $Y_{i}$, for each $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite. Let $P_{1},\ldots,P_{r}$ be vertex-disjoint paths satisfying (P’1)(P’3), guaranteed by Lemma 7.2, and let $Q\coloneqq P_{1}\cup\ldots\cup P_{r}$.
@@ 7.3.1 Constructing a balanced pair of $F$-subdivisions in each expander
Thus, by applying (ii) of Lemma 7.1 (with $2\xi$ playing the role of $\xi$), we obtain an $(s,s^{\prime})$-path $P$ in $G^{\prime}[Z_{i}]-W$ of length at most $15/\delta$. But then $\mathcal{P}\cup\{P\}$ contradicts the maximality of $\mathcal{P}$. This shows that $|\mathcal{P}|=2e(F)$, as claimed.
Pick a collection $\mathcal{P}$ as guaranteed by the previous paragraph. Take $F_{i}$ to be the subgraph consisting of the paths in $\mathcal{P}$ with ends in $S$, and take $F_{i}^{\prime}$ to be the subgraph consisting of the paths in $\mathcal{P}$ with ends in $S^{\prime}$. We claim that $F_{i}$ and $F_{i}^{\prime}$ satisfy the requirements of Claim 7.3. Indeed, first note that $F_{i}$ is an $F$-subdivision in $G^{\prime}[Z_{i}]-V(Q)$ whose branch vertices are in $S$, and similarly $F_{i}^{\prime}$ is an $F$-subdivision in $G^{\prime}[Z_{i}]-V(Q)$ whose branch vertices are in $S^{\prime}$. Second, notice that $F_{i}$ and $F_{i}^{\prime}$ are vertex-disjoint, since the paths in $\mathcal{P}$ are internally vertex-disjoint whose internal vertices are in $Z_{i}-(V(Q)\cup S\cup S^{\prime})$. Third, we have $|V(F_{i})|,|V(F_{i}^{\prime})|\leq e(F)\cdot 15/\delta$, since every path in $\mathcal{P}$ has length at most $15/\delta$ by our choice of $\mathcal{P}$. Finally, if $G[Z_{i}]$ is $\beta$-almost-bipartite, then for every path $P\in\mathcal{P}$ with ends in $S$, we have
@@
This proves (23), completing the proof of Claim 7.3. ∎
@@ 7.3.2 Absorbing the paths $P_{1},\ldots,P_{r}$ and all the uncovered vertices into the subdivisions $F_{i}$
For every $i\in[r]$, let $F_{i},F^{\prime}_{i}$ be the two vertex-disjoint subdivisions of $F$ in $G^{\prime}[Z_{i}]-V(Q)$, as guaranteed by Claim 7.3. Denote the leaves of $P_{i}$ by $u_{i}$ and $v_{i}$, where $u_{i}\in X_{i}$ and $v_{i}\in Y_{i}$ if $G[Z_{i}]$ is $\beta$-almost-bipartite. Let $x_{i}y_{i}$ be an arbitrary edge of $F_{i}$, where $x_{i}\in X_{i}$ and $y_{i}\in Y_{i}$ if $G[Z_{i}]$ is $\beta$-almost-bipartite. Let $Q_{i}$ be a shortest $(y_{i},u_{i})$-path in $G^{\prime}[Z_{i}]$ whose interior avoids $V(Q)\cup V(F_{i})\cup V(F_{i}^{\prime})$. By (P’1),
@@
Hence, it follows that $Q_{i}$ exists and is of length at most $15/\delta$ by applying (ii) of Lemma 7.1 (with $\sqrt{\xi}$ playing the role of $\xi$). Let $H\coloneqq G[Z_{i}]$, and let $H^{\prime}\coloneqq G^{\prime}[Z_{i}]$. Then, by Lemma 7.1, $H^{\prime}$ has no $\zeta/2$-sparse cuts. Now let
@@
showing that $H^{\prime}$ has average degree at least $d-(2\beta/\delta^{2})|V(H^{\prime})|$. Moreover, $H^{\prime}$ is either bipartite or $\gamma$-far-from-bipartite by definition. Therefore, using $|W_{i}|\leq 2\sqrt{\xi}n$, we can apply Lemma 3.3 with $H^{\prime}$, $W_{i}$, $2\beta/\delta^{2}$, $2\xi$, $\zeta/2$ playing the roles of $G$, $W$, $\eta$, $\xi$, $\zeta$, respectively, to find a Hamilton path $Q_{i}^{\prime}$ in $H^{\prime}-W_{i}=G^{\prime}[Z_{i}]-W_{i}$ with ends $x_{i}$ and $v_{i}$. Note that Lemma 3.3 indeed applies because when $G[Z_{i}]$ is $\beta$-almost-bipartite (so $H^{\prime}$ is bipartite), (P’3) and (23) together imply that $|X_{i}-W_{i}|=|Y_{i}-W_{i}|$.